An automobile is moving down a straight road and braked with a constant acceleration of -5 m/s2. If the initial velocity of the car is 35 m/s and the final velocity of the car is 15 m/s, calculate the displace
1
Expert's answer
2020-12-06T17:26:19-0500

The velocity "v" of an object moving with a constant acceleration at any moment of time "t" is given by the following formula:


"v = v_0+at"


where "v_0 =35m\/s" is the initial velocity of the car and "a = -5m\/s^2" is its acceleration. If we set "v = 15m\/s", we can find the time that was taken to reach from "v_0" to "v":


"t = \dfrac{v-v_0}{a} = \dfrac{15-35}{-5} = 4s"

The displacement after this time is:


"d = v_0t + \dfrac{at^2}{2}\\
d = 35\cdot 4 - \dfrac{5\cdot 4^2}{2} = 100m"

Answer. 100 m.